OpenStudy (anonymous):

Blocks numbered 0 through 9 are placed in a box, and a block is randomly picked. The probability of picking an odd prime number is 1/5 I THINK The probability of picking a number greater than 0 that is also a perfect square is BLANK? IDK THAT ONE

OpenStudy (anonymous):

whats a perfect square also

OpenStudy (anonymous):

oh a odd prime number

OpenStudy (anonymous):

whats a prime number

jimthompson5910 (jim_thompson5910):

A prime number is a number that only has factors of 1 and itself example: 29 is a prime number since only 1 and 29 are factors of 29

OpenStudy (anonymous):

1/3 is the odd prime number answer

jimthompson5910 (jim_thompson5910):

btw, the number 1 is NOT prime

OpenStudy (anonymous):

I THINK

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

im staying with 1/3 now i only dont understand what a perfect square is

jimthompson5910 (jim_thompson5910):

a perfect square is a whole number that is the result of squaring some smaller whole number example: 49 is a perfect square because 7^2 = 7*7 = 49

OpenStudy (anonymous):

alright so

jimthompson5910 (jim_thompson5910):

here are all the blocks laid out 

jimthompson5910 (jim_thompson5910):

circle the blocks that are prime numbers click the pencil of my drawing so you can draw on it

OpenStudy (anonymous):